40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 17 12:02:55 2014, Koji, Summary, General, IFO recovery / confirmed ETMX in trouble MCspots.png
    Reply  Mon Mar 17 20:14:34 2014, Koji, Summary, General, IFO recovery / confirmed ETMX in trouble ETMX_UL.JPGETMX_UR.JPGETMX_LR.JPGETMX_LL.JPG
Message ID: 9731     Entry time: Mon Mar 17 12:02:55 2014     Reply to this: 9733
Author: Koji 
Type: Summary 
Category: General 
Subject: IFO recovery / confirmed ETMX in trouble 

I confirmed that we need to vent the chambers.

All of the mirrors have been aligned except for ETMX.

ETMX does not respond to the excitation by the UR and LR coils. Likely that the magnets are knocked off, or stuck in the coil.

PRM/SRM oplevs are too much off and can't be turned on. Need realignment of the beams on the QPDs.


- FB was down. FB restarted ("telnet fb 8087", then type shutdown)

- Aligned the MC mirrors.

- Aligned PRM. Look at the REFL. It was slightly mislisligned.

- AS has no beam. The Y arm was resonating with the green. So I determined that the TTs were the misaligned guys.

- Touched TT2 pitch with an increment of 0.1. Immediately the AS beam spot for ITMY was found. And the arm was resonating.

- The RM was further aligned. The bias sliders were saved and then the PRM was misaligned.

- Yarm was locked on TEM01. The ASS maximized the transmission for TEM01, and then the arm was locked on TEM00.
  The ASS aligned the arm and TTs. These values were saved.

- Yarm was aligned and I can see the AS spot. So I believe the BS is still well aligned.

- Aligned the PRM to reduce the ghost beams.

- Moved the ITMX to have Michelson fringes properly.

- Also aligned the SRM.

- Now ETMX was checked. Played with the alignment biases to see if the mirror was sticking on the coils. The mirror can rock a little, but it did not come back.

- Then, checked each magnets. 0.8Hz 1000cnt signals were injected to each coils (cf. C1:SUS-PRM_**COIL_EXC) to see how the mirror could react.
  The OSEM output and green spot on the ETMX cage were observed.

- Saw some response by actuating the UL, LL, SD coils.

- Saw no response from the UR and LR coils. They show the OSEM output of zero. Does this mean the magnets fell in the coils?

//Manasa// MC spot positions measured and they look alright with not much change from before the earthquake (attach)

Attachment 1: MCspots.png  95 kB  Uploaded Mon Mar 17 13:54:55 2014  | Hide | Hide all
MCspots.png
ELOG V3.1.3-